Suppose that a 15-year $1000-face value government (risk-free) coupon bond with 8% coupon payments (paid annually) is currently selling at par. ABC Company’s bonds are rated AA, and the credit spread on these bonds is 3%. At what price would you be willing to purchase a 15-year ABC Company bond with a $1000-face value if it offers 8% coupon payments (paid annually)? Show your calculations and explain your result.
Answer #1

As risk free bonds are selling at par, yield on risk free bonds=coupon rate=8%

yield on ABC bonds=risk free bond yield+spread=8%+3%=11%

Price of bond=Face Value*Coupon Rate/yield*(1-1/(1+yield)^t)+Face Value/(1+yield)^t=1000*8%/11%*(1-1/1.11^15)+1000/1.11^15=784.27391

One would be willing to purchase the bond if market price is less than 784.27391